# SelfDefend: LLMs Can Defend Themselves against Jailbreaking in a Practical Manner

### v1

@@ 5.3 Measurement Results
Table III summarizes our core measurement results, which we separate into the defense effectiveness against jailbreaks and the impact on normal user prompts. Additionally, we measure the extra delay introduced by SelfDefend.
@@
Baseline Performance. We first measure the baseline performance of using GPT-3.5/4 only under the tested jailbreaks. As shown in Table III, these jailbreaks are much more effective on GPT-3.5 compared to GPT-4. The ASR for GPT-3.5 can be as high as 0.39 to 0.98 (average: 0.75), whereas the ASR for GPT-4 typically ranges from 0.08 to 0.33, except for indirect jailbreaks in which case GPT-4 also has a high ASR (0.74 for DrAttack and 0.9 for Puzzler).
Defense Effectiveness. We then compare ASR after equipping SelfDefend with the baseline model. In ablation, we also report ASR using only the shadow stack. Our findings are summarized as follows:
  • Based on GPT-3.5, although the SelfDefend frameworks supported by intent prompt and direct prompt both decrease success rates, there is a considerable difference in the enhancement of defense performance achieved by the two methods. The SelfDefend framework with direct prompt shows a modest reduction in terms of ASR due to its internal shadow stack performing even worse than GPT-3.5. For example, its ASR on AutoDAN is 96%, compared to GPT-3.5’s 90%. In contrast, the SelfDefend framework utilizing intent prompt demonstrates significantly better performance than the one with direct prompt, with the reduction factor ranging from as high as 95.74% to as low as 8.97%.
  • Based on GPT-4, both SelfDefend frameworks significantly reduce the ASRs of various jailbreak attacks. Different from the results on GPT-3.5, SelfDefend with direct prompt is better than SelfDefend with intent prompt due to the difference between the two shadow stacks. Although jailbreak attacks present lower ASRs than on GPT-3.5, the reduction factor of SelfDefend with direct prompt ranges from 36.36% to 100.00%.
  • SelfDefend with direct prompt and SelfDefend with intent prompt present the best defense enhance on GPT-4 and GPT-3.5, respectively. Furthermore, SelfDefend with direct prompt demonstrates the higher reduction factor than SelfDefend with intent prompt on GPT-3.5. Thus, we contend that SelfDefend with direct prompt relies more on the model’s own discrimination ability so that is suitable for powerful models such as GPT-4, while the intent prompt can strengthen the model’s logical reasoning and still has decent defense capabilities even on weaker models.
@@
Impact on Normal Queries. Besides defense effectiveness, we also measure the impact of SelfDefend on normal queries by evaluating the ASR of 805 normal prompts from AlpacaEval. Table III shows that for both GPT-3.5 and GPT-4, SelfDefend (with $P_{direct}$) slightly reduces the ASR by 2% (from 0.97 to 0.95) and 3% (from 0.97 to 0.94), respectively. Meanwhile, SelfDefend (with $P_{direct}$) does not decrease the ASR at all for either model. Considering that the ASR under SelfDefend is still very high and close to the baseline, we conclude that SelfDefend’s defense would not notably impact normal user queries.
@@
Extra Delay $\Delta d$. Recall that one of our design objectives is to incur negligible delay to user prompts (O2 in §3.1). We thus further measure the extra delay $\Delta d$ caused by SelfDefend, which is calculated as follows:
@@
where $d_{total}$ denotes the total delay to SelfDefend-protected LLMs, and $d_{normal}$ represents the separate time cost of target LLMs in the normal stack. The metric of $\Delta d$ is mainly important for normal prompts because it would directly affect the experience of normal users. Figure 3(a) plots the cumulative distribution function (CDF) of $\Delta d$ for normal prompts. We can see that over 95% of the cases incur zero delay across three configurations of SelfDefend, demonstrating that SelfDefend is capable of defending against jailbreaks at the cost of negligible delay for normal user prompts. The only exception occurs with SelfDefend using $P_{intent}$ and GPT-3.5, where around 80% of the cases have no extra delay.
For jailbreak prompts, we also calculate their average $\Delta d$ for different attack scenarios in Figure 3(b). We can see that GPT-3.5’s average $\Delta d$ is under 0.1 seconds for all eight scenarios, while GPT-4 incurs negligible $\Delta d$ only in the normal scenario of AlpacaEval. This is likely because the baseline GPT-4 has good defense performance for most jailbreaks, suggesting its original output is already short and the delay is minimal.
Figure 4: The training procedure for fine-tuning our defense model.
@@ 6 Tuning a Dedicated Defense Model
Based on the empirical measurement results, we have observed that GPT-4, when equipped with SelfDefend, significantly reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.06 (for GPT-4-based SelfDefend with $P_{direct}$). However, GPT-4 is commonly known to be expensive [58], and its closed-source nature raises privacy concerns for non-OpenAI model vendors. Therefore, our objective in this section is to tune an open-source model that can be used under the SelfDefend architecture for robust, low-cost, and self-contained jailbreak defense.
@@ 6.1 Design Overview
Given the powerful defensive capability of GPT-4-based SelfDefend, our intuition is to “transfer” this capability to an open-source model. To do so, we leverage GPT-4-based SelfDefend to distill and generate high-quality tuning data. Figure 4 depicts our proposed pipeline to fine-tune an open-source defense model. Specifically, by continuously incorporating harmful and harmless prompts into our defense prompts (i.e., $P_{direct}$ or $P_{intent}$) as inputs for GPT-4, we gather their outputs as labels for these samples. Since we utilize two defense prompts, we eventually obtain two separate datasets, which we then use to fine-tune the employed open-source model for our defense purposes.
@@ 6.2 Data Distillation
To distill GPT-4’s “knowledge” on safety policies, a dataset containing both harmful and harmless queries is necessary. We utilize the red-team data from Anthropic [59] as the query dataset. This dataset consists of 38,961 text transcripts documenting conversations between human adversaries and AI assistants. From this dataset, we select the initial human prompt and exclude the corresponding assistant response, omitting any subsequent exchanges, to create a single-turn prompt dataset designated as $\mathcal{D}_{red}$. It is worth noting that this dataset is released in 2022 and has no overlap with JailbreakBench/MultiJail released in 2024. Furthermore, this dataset is exclusively in English, implying that a model trained on this data will not acquire any additional multilingual capabilities.
@@ 6.3 LoRA Fine-Tuning
We allocate 80% of the samples in $\mathcal{D}_{dir}$ or $\mathcal{D}_{int}$ for fine-tuning and reserve the remaining 20% as validation sets. During the training process, we continue to use our defense prompts and queries from the collected datasets as inputs; that is, we take $P_{dir}(x)$ or $P_{int}(x)$ as input and $y$ as the label to fine-tune the Llama model.
@@
where $\Theta$ represents the trainable parameters of LoRA, $\Phi_{0}$ denotes the pre-trained weights of the Llama model, and $\Delta\Phi(\Theta)$ indicates the parameter increment determined by LoRA. Fine-tuning the Llama model with this objective can enhance its ability to directly detect jailbreak portions in query prompts.
@@
This objective allows the fine-tuned model to identify jailbreak intentions within query prompts.
Implementation. We use LoRA to fine tune the Llama model with $r$ of 8 and $\alpha$ of 32 [60]. The initial learning rate is $10^{-3}$. The training epochs are 1 in batch size of 8.
@@ 7 Evaluation
TABLE IV: Jailbreak ASR for various defense methods.
Target Model Defense Method Optimization-based Template-based Indirect Multilingual Normal
GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 GPT-3.5 (baseline) 0.56 0.94 0.72 0.90 0.78 0.98 0.39 0.97
Perplexity Filter [62] 0.03 1.00 1.00 1.00 1.00 1.00 1.00 0.99
SmoothLLM [51] 0.48 0.89 0.65 0.93 0.85 0.99 1.00 0.96
Llama Guard [42] 0.41 0.61 0.43 1.00 0.97 0.93 0.95 0.99
Llama Guard 2 [63] 0.14 0.50 0.37 0.15 0.89 0.64 0.55 0.99
$P_{direct}$-tuned Shadow Stack 0.08 0.11 0.14 0.07 0.78 0.07 0.74 0.96
$P_{direct}$-tuned SelfDefend 0.06 0.11 0.07 0.07 0.62 0.07 0.30 0.94
$P_{intent}$-tuned Shadow Stack 0.08 0.05 0.16 0.09 0.18 0.20 0.57 0.99
$P_{intent}$-tuned SelfDefend 0.05 0.05 0.12 0.08 0.16 0.20 0.26 0.97
Double shadow stack 0.04 0.02 0.10 0.05 0.18 0.07 0.47 0.97
Double shadow stack + GPT-3.5 0.03 0.02 0.06 0.05 0.16 0.07 0.19 0.95
GPT-4 GPT-4 (baseline) 0.08 0.16 0.33 0.19 0.74 0.90 0.08 0.97
Perplexity Filter [62] 0.03 1.00 1.00 1.00 1.00 1.00 1.00 0.99
SmoothLLM [51] 0.07 0.19 0.33 0.18 0.91 0.88 1.00 0.97
Llama Guard [42] 0.41 0.61 0.43 1.00 0.98 0.93 0.95 0.99
Llama Guard 2 [63] 0.14 0.50 0.35 0.15 0.91 0.64 0.55 0.99
$P_{direct}$-tuned Shadow Stack 0.04 0.11 0.17 0.07 0.71 0.12 0.71 0.96
$P_{direct}$-tuned SelfDefend 0.01 0.07 0.15 0.05 0.58 0.07 0.06 0.94
$P_{intent}$-tuned Shadow Stack 0.08 0.03 0.19 0.07 0.18 0.19 0.56 0.99
$P_{intent}$-tuned SelfDefend 0.04 0.02 0.19 0.06 0.17 0.14 0.04 0.97
Double shadow stack 0.02 0.02 0.12 0.04 0.18 0.12 0.47 0.97
Double shadow stack + GPT-4 0.01 0.02 0.12 0.04 0.17 0.07 0.04 0.95
Llama 2 Llama-2-7b-chat (baseline) 0.57 0.96 0.59 0.68 0.88 0.99 0.14 0.98
Perplexity Filter [62] 0.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99
SmoothLLM [51] 0.93 1.00 0.82 0.95 0.98 1.00 0.99 0.99
Llama Guard [42] 0.40 0.41 0.46 1.00 0.84 0.93 0.95 0.99
Llama Guard 2 [63] 0.17 0.53 0.41 0.15 0.89 0.64 0.55 0.99
(7b-chat) $P_{direct}$-tuned Shadow Stack 0.06 0.06 0.25 0.05 0.36 0.09 0.73 0.97
$P_{direct}$-tuned SelfDefend 0.04 0.06 0.22 0.05 0.31 0.09 0.10 0.95
$P_{intent}$-tuned Shadow Stack 0.11 0.02 0.24 0.08 0.15 0.22 0.58 0.99
$P_{intent}$-tuned SelfDefend 0.09 0.02 0.21 0.07 0.15 0.22 0.06 0.98
Double shadow stack 0.05 0.02 0.18 0.04 0.12 0.04 0.48 0.97
Double shadow stack + Llama 2 0.03 0.02 0.16 0.04 0.11 0.04 0.06 0.95
In this section, we extensively evaluate SelfDefend with Llama-2-tuned defense models and compare their performances with other jailbreak defense methods. We generally follow the setup mentioned in §5 unless explicitly specified in §7.1. For each evaluation target, we assess not only the defense effectiveness (in §7.2) and the extra delay $\Delta d$ (in §7.3) as in §5, but also whether the detected harmful portion actually aligns with the original prompt (i.e., explainability in §7.4), and how robust our defense models are against targeted optimization-based attacks (in §7.5) and prompt injection (in §7.6).
@@ 7.1 Experimental Setup
Target LLMs, Benchmarks, and Environment. As shown in Table IV, our tuned defense models are paired with protect three popular LLMs: GPT-3.5, GPT-4, and the open-source Llama-2-7b-chat [61]. We use the same benchmark datasets as in Table II (see §5). We follow the same procedure and configuration as in §5 to generate the tested jailbreak prompts for GPT-3.5 and GPT-4. The process is similar for testing the target model Llama-2-7b-chat; we simply set the target model for query-based JSAA, PAIR, and DrAttack jailbreaks to Llama-2-7b-chat to generate their jailbreak prompts. Our evaluation is implemented using PyTorch 2.2.1 and conducted on an NVIDIA TESLA H800 GPU.
Baselines. We compare our framework with popular jailbreak defense methods, including Perplexity Filter [62], SmoothLLM [51], Llama Guard [42], and Llama Guard 2 [63]. Specifically, Perplexity Filter leverages a Llama-2-7b model to calculate the perplexity of the input prompt. A jailbreak is considered to happen when the perplexity exceeds a threshold. We set this threshold at the maximum perplexity of any prompt in the JailbreakBench dataset of harmful behavior prompts. SmoothLLM perturbs the jailbreak prompts with character-level changes to enable the target LLM to perform defense. In this paper, we set SmoothLLM to perform character swapping with a 10% perturbation percentage. Llama Guard is a fine-tuned Llama-2-7b model designed to detect the toxicity category of input prompts. Llama Guard 2 is a fine-tuned version of Llama-3-8B [81].
Metric. We measure the performance of defense methods by the attack success rate (ASR), i.e., the frequency with which jailbreak prompts in a benchmark dataset bypass the detection of a defense method. The lower the ASR, the stronger the defense performance. However, the definition of “attack success” is different for different defense methods. For Perplexity Filter, the success denotes that the perplexity of the query text is no more than the fixed threshold. For SmoothLLM, the definition of success is the success of the attack method, i.e., ASR [29] or Unsafe Rate [34]. It is worth noting that the Unsafe Rate refers to the unsafety of the target model’s response, rather than the input prompt. For Llama Guard, it is success when its response is “unsafe”. For our fine-tuned models, the success represent their response being “No.”
Figure 5: An example of tuned models against GCG on GPT-3.5.
Figure 6: An example of tuned models against PAIR on GPT-3.5.
Figure 7: An example of tuned models against DrAttack on GPT-3.5.
Figure 8: An example of tuned models against MultiJail on GPT-3.5.
Figure 9: Tuned models’ response to a normal prompt from AlpacaEval.

### v2

@@ 5.3 Measurement Results
Table 3 summarizes our core measurement results, which are divided into two categories: the defense effectiveness against jailbreaks and the impact on normal user prompts. Additionally, we measure the extra delay introduced by SelfDefend.
@@
Baseline Performance. We first measure the baseline performance of using GPT-3.5/4 only under the tested jailbreaks. As shown in Table 3, these jailbreaks are much more effective on GPT-3.5 compared to GPT-4. The ASR for GPT-3.5 can be as high as 0.256 to 0.980 (average: 0.657), whereas the ASR for GPT-4 typically ranges from 0.047 to 0.330, except for indirect jailbreaks where GPT-4 also exhibits a high ASR (0.74 for DrAttack and 0.900 for Puzzler).
Defense Effectiveness. We then compare ASR after equipping the baseline model with SelfDefend and report additional ASR findings when using only the shadow stack for ablation exploration. For GPT-3.5, both SelfDefend versions—supported by $P_{direct}$ and $P_{intent}$—reduce ASRs but show varying degrees of defense enhancement. SelfDefend with $P_{direct}$ exhibits a modest reduction, with its shadow stack performing worse than GPT-3.5 itself; for instance, its ASR on AutoDAN is 96%, compared to GPT-3.5’s 90%. In contrast, SelfDefend with $P_{intent}$ significantly outperforms the $P_{direct}$ version, achieving a reduction factor ranging from 97.26% to 8.97%.
For GPT-4, both SelfDefend versions substantially lower the ASRs across various jailbreak attacks. Unlike with GPT-3.5, SelfDefend with $P_{direct}$ outperforms $P_{intent}$ due to differences in shadow stack effectiveness, with reduction factors ranging from 69.69% to 100.00%. SelfDefend with $P_{direct}$ provides the best defense enhancement on GPT-4, while SelfDefend with $P_{intent}$ excels on GPT-3.5. SelfDefend with $P_{direct}$ relies more on the model’s inherent discrimination capabilities, making it more suitable for advanced models like GPT-4. Conversely, the intent prompt enhances logical reasoning and remains effective even on less advanced models.
@@
Impact on Normal Queries. Besides defense effectiveness, we also measure the impact of SelfDefend on normal queries by evaluating the ASR of 805 normal prompts from AlpacaEval. Table 3 shows that for both GPT-3.5 and GPT-4, SelfDefend (with $P_{direct}$) slightly reduces the ASR by 2% (from 0.977 to 0.957) and 2.7% (from 0.973 to 0.946), respectively. Meanwhile, SelfDefend (with $P_{intent}$) hardly decreases the ASR for either model. Considering that the ASR under SelfDefend is still very high and close to the baseline, we conclude that SelfDefend’s defense would not notably impact normal user queries.
@@
Extra Delay $\Delta d$. Recall that one of our design objectives is to incur negligible delays to user prompts (O2 in §3.1). We thus further measure the extra delay $\Delta d$ caused by SelfDefend, which is calculated as follows:
@@
where $d_{total}$ denotes the total delay to SelfDefend-protected LLMs, and $d_{normal}$ represents the separate time cost of the target LLM in the normal stack. The metric of $\Delta d$ is essential for normal prompts because it would directly affect the experience of normal users. Figure 3(a) plots the cumulative distribution function (CDF) of $\Delta d$ for normal prompts. We can see that over 95% of the cases incur zero delays across three configurations of SelfDefend, demonstrating that SelfDefend is capable of defending against jailbreaks at the cost of negligible delay for normal users. The only exception occurs with SelfDefend using $P_{intent}$ and GPT-3.5, where around 80% of the cases have no extra delay.
For jailbreak prompts, we also calculate their average $\Delta d$ for different attack scenarios in Figure 3(b). We observe that GPT-3.5’s average $\Delta d$ is under 0.1 seconds for all eight scenarios, while GPT-4 incurs negligible $\Delta d$ only in the normal scenario of AlpacaEval. This is likely because the baseline GPT-4 has good defense performance for most jailbreaks, with its original output already short and the delay minimal.
Figure 4: The training procedure for fine-tuning our open-source defense model.
@@ 6 Tuning a Dedicated Defense Model
Based on the empirical measurement results, we have observed that GPT-4, when equipped with SelfDefend, significantly reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.063. However, GPT-4 is known to be expensive [3], and its closed-source nature raises privacy concerns for non-OpenAI model vendors. Therefore, our objective in this section is to tune an open-source model that can be used under the SelfDefend architecture for robust, low-cost, and self-contained defense.
@@ 6.1 Design Overview
Given the powerful defensive capability of GPT-4-based SelfDefend, our intuition is to “transfer” this capability to an open-source model. To do so, we leverage GPT-4-based SelfDefend to distill and generate high-quality tuning data. Figure 4 depicts our pipeline to fine-tune an open-source defense model. Specifically, by continuously incorporating harmful and harmless prompts into our defense prompts (i.e., $P_{direct}$ or $P_{intent}$) as inputs for GPT-4, we gather their outputs as labels for these samples. Since we utilize two defense prompts, we eventually obtain two separate datasets, which we then use to fine-tune the employed open-source model.
@@ 6.2 Data Distillation
To distill GPT-4’s “knowledge” on safety policies, a dataset containing both harmful and harmless queries is necessary. We utilize the red-team data from Anthropic [19] as the query dataset. This dataset consists of 38,961 text transcripts documenting conversations between human adversaries and AI assistants. From this dataset, we select the initial human prompt and exclude the corresponding assistant response, omitting any subsequent exchanges, to create a single-turn prompt dataset designated as $\mathcal{D}_{red}$. Note that this dataset was released in 2022 and has no overlap with JailbreakHub, JailbreakBench, or MultiJail, which were released in 2023 or 2024. Furthermore, this dataset is exclusively in English, implying that a model trained on this data will not acquire any additional multilingual capabilities.
@@ 6.3 LoRA Fine-Tuning
where $\Theta$ represents the trainable parameters of LoRA, $\Phi_{0}$ denotes the pre-trained weights of the Llama model, and $\Delta\Phi(\Theta)$ indicates the parameter increment determined by LoRA. Fine-tuning the Llama model with this objective can enhance its ability to directly detect jailbreak parts in query prompts.
@@
Implementation. We use LoRA to fine tune the Llama model with $r$ of 8 and $\alpha$ of 32 [22]. The initial learning rate is $10^{-3}$. The training epochs are 1 in a batch size of 8. We allocate 80% of the samples in $\mathcal{D}_{dir}$ or $\mathcal{D}_{int}$ for fine-tuning and reserve the remaining 20% as validation sets. During the training process, we continue to use our defense prompts and queries from the collected datasets as inputs; that is, we take $P_{dir}[x]$ or $P_{int}[x]$ as input and $y$ as the label to fine-tune the Llama model.
Table 4: Jailbreak ASR for various defense methods. For ICD, SmoothLLM and SafeDecoding, we present the performance of their enhanced models. For detection-based Perplexity Filter and Llama Guard (2), we report ASRs only on their detection modules. Since SafeDecoding works for a white-box target model, we show its results only on the publicly available Llama-2.
Target Defense Method Human Optimization Generation Indirect Multilingual Normal
DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 GPT-3.5 (baseline) 0.256 0.560 0.900 0.720 0.670 0.780 0.980 0.393 0.977
ICD [73] 0.226 0.230 0.840 0.360 0.330 0.750 0.990 0.321 0.960
Perplexity Filter [26] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [59] 0.238 0.480 0.930 0.650 0.610 0.850 0.990 0.705 0.968
Llama Guard [25] 0.561 0.410 0.580 0.430 0.470 0.970 0.930 0.952 0.996
Llama Guard 2 [66] 0.441 0.140 0.150 0.370 0.360 0.890 0.640 0.559 0.991
$P_{direct}$-tuned Shadow Stack 0.262 0.080 0.070 0.140 0.210 0.780 0.070 0.749 0.968
$P_{direct}$-tuned SelfDefend 0.111 0.060 0.070 0.070 0.170 0.620 0.070 0.302 0.948
$P_{intent}$-tuned Shadow Stack 0.297 0.080 0.090 0.160 0.240 0.180 0.200 0.578 0.996
$P_{intent}$-tuned SelfDefend 0.125 0.050 0.080 0.120 0.200 0.160 0.200 0.260 0.975
Double shadow stack 0.213 0.040 0.050 0.100 0.130 0.180 0.070 0.470 0.966
Double shadow stack+GPT-3.5 0.091 0.030 0.050 0.060 0.100 0.160 0.070 0.187 0.947
GPT-4 GPT-4 (baseline) 0.047 0.080 0.190 0.330 0.310 0.740 0.900 0.076 0.973
ICD [73] 0.062 0.050 0.030 0.230 0.230 0.430 0.640 0.051 0.970
Perplexity Filter [26] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [59] 0.030 0.070 0.180 0.330 0.330 0.910 0.880 0.651 0.971
Llama Guard [25] 0.561 0.410 0.580 0.430 0.400 0.980 0.930 0.952 0.996
Llama Guard 2 [66] 0.441 0.140 0.150 0.350 0.330 0.910 0.640 0.559 0.991
$P_{direct}$-tuned Shadow Stack 0.262 0.040 0.070 0.170 0.220 0.710 0.120 0.717 0.969
$P_{direct}$-tuned SelfDefend 0.032 0.010 0.050 0.150 0.150 0.580 0.070 0.060 0.947
$P_{intent}$-tuned Shadow Stack 0.284 0.080 0.070 0.190 0.190 0.180 0.190 0.565 0.995
$P_{intent}$-tuned SelfDefend 0.034 0.040 0.060 0.190 0.160 0.170 0.140 0.044 0.970
Double shadow stack 0.198 0.020 0.040 0.120 0.130 0.180 0.120 0.467 0.968
Double shadow stack+GPT-4 0.026 0.010 0.040 0.120 0.110 0.170 0.070 0.038 0.945
Llama-2 Llama-2-7b-chat (baseline) 0.678 0.570 0.680 0.590 0.610 0.880 0.990 0.143 0.988
ICD [73] 0.474 0.700 0.560 0.310 0.320 0.220 1.000 0.946 0.898
SafeDecoding [82] 0.655 0.550 0.740 0.560 0.590 0.640 1.000 0.857 0.981
Perplexity Filter [26] 1.000 0.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [59] 0.681 0.930 0.950 0.820 0.810 0.980 1.000 1.000 0.993
Llama Guard [25] 0.561 0.400 0.580 0.460 0.420 0.840 0.930 0.952 0.996
Llama Guard 2 [66] 0.441 0.170 0.150 0.410 0.350 0.890 0.640 0.559 0.991
(7b-chat) $P_{direct}$-tuned Shadow Stack 0.257 0.060 0.050 0.250 0.190 0.360 0.090 0.737 0.970
$P_{direct}$-tuned SelfDefend 0.214 0.040 0.050 0.220 0.180 0.310 0.090 0.102 0.959
$P_{intent}$-tuned Shadow Stack 0.289 0.110 0.080 0.240 0.140 0.150 0.220 0.587 0.991
$P_{intent}$-tuned SelfDefend 0.242 0.090 0.070 0.210 0.140 0.150 0.220 0.063 0.980
Double shadow stack 0.198 0.050 0.040 0.180 0.140 0.120 0.040 0.479 0.965
Double shadow stack+Llama-2 0.164 0.030 0.040 0.160 0.140 0.110 0.040 0.057 0.954
@@ 7 Evaluation
In this section, we extensively evaluate SelfDefend with Llama-2-tuned defense models and compare their performance with other jailbreak defense methods. We generally follow the setup mentioned in §5 unless explicitly specified in §7.1. For each evaluation target, we assess not only the defense effectiveness (in §7.2) and the extra delay $\Delta d$ (in §7.3) as in §5, but also whether the detected harmful portion actually aligns with the original prompt (i.e., explainability in §7.4), and how robust our defense models are against adaptive attacks (in §7.5) and prompt injection (in §7.6).
@@ 7.1 Experimental Setup
Target LLMs, Benchmarks, and Environment. As shown in Table 4, our tuned defense models are paired with protect three popular LLMs: GPT-3.5, GPT-4, and the open-source Llama-2-7b-chat [67]. We use the same benchmark datasets as in Table 2 (see §5). We follow the same procedure and configuration as in §5 to generate the tested jailbreak prompts for GPT-3.5 and GPT-4. The process is similar for testing the target model Llama-2-7b-chat; we set the target model for query-based PAIR, TAP and DrAttack to Llama-2-7b-chat to generate their jailbreak prompts. Our evaluations are implemented using PyTorch 2.2.1 and conducted on an NVIDIA TESLA H800 GPU.
Baselines. We compare our framework with popular jailbreak defense methods, including ICD [73], SafeDecoding[82], Perplexity Filter [26], SmoothLLM [59], Llama Guard [25] and Llama Guard 2 [66]. Specifically, ICD adds in-context demonstrations in input prompts to enhance the safety of the target model. We adopt the same 1-shot demonstration as the ICD in [95]. SafeDecoding seeks to methodically scrutinize safety-related disclaimers and amplify the probabilities of their associated token sequences. We only show the defense performance of SafeDecoding on Llama-2-7b-chat since it requires fine-tuning an expert model based on the target model. Perplexity Filter leverages a Llama-2-7b model to calculate the perplexity of the input prompt. A jailbreak is considered to happen when the perplexity exceeds a threshold. We set this threshold at the maximum perplexity of any prompt in the JailbreakBench dataset of harmful behavior prompts. SmoothLLM perturbs the jailbreak prompts with character-level changes to enable the target LLM to perform defense. In this paper, we set SmoothLLM to conduct character swapping with a 10% perturbation percentage. Llama Guard is a fine-tuned Llama-2-7b model designed to detect the toxicity category of input prompts. Llama Guard 2 is a fine-tuned version of Llama-3-8B [18].
Metric. We measure the performance of defense methods by the attack success rate (ASR), i.e., the frequency with which jailbreak prompts in a benchmark dataset bypass the guardrail of a defense method. The lower the ASR, the stronger the defense performance. However, the definition of “attack success” is different for distinct defense methods. For ICD and SafeDecoding, we adopt the above ASR [99] or Unsafe Rate [17] to measure their defense capability since they directly enhance the model safety instead of plug-in jailbreak detection. For Perplexity Filter, the success denotes that the perplexity of the query text is no more than the fixed threshold. For SmoothLLM, the definition is the success of the attack method, i.e., ASR [99] or Unsafe Rate [17]. It is worth noting that the Unsafe Rate refers to the unsafety of the target model’s response, rather than the input prompt. For Llama Guard, it is a success when its response is “unsafe”. For our fine-tuned models, the success represents their response being “No”.
Figure 5: Responses of defense models against a jailbreak prompt of AutoDAN.